A researcher found a 95% confidence interval for the proportion of college students who believe in ghosts.

The interval was (0.6525, 0.7535). What is the margin of error for this estimate?
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.6525+0.7535)./ 2=0.703

Margin of error = E = Upper confidence interval - = 0.7535-0.703=0.0505